OpenStudy (anonymous):

Last one!!! What is the y-intercept of a line through (2,7) and perpendicular to the graph of y=-3/2x+6???

OpenStudy (anonymous):

with this be a form of a linear equation ?

OpenStudy (anonymous):

yes

OpenStudy (austinl):

What you would want to do here is have another line in the form of slope-intercept. \(\displaystyle y=mx+b\) Now for a line to be perpendicular, that would mean that the slope would be the negative reciprocal of the given slope. \(\displaystyle m_{given}=-\frac{3}{2},~m_{perpendicular}=\frac{2}{3}\) To give us the very basic equation, \(\displaystyle y=\frac{2}{3}x+b\) Now to get the y-intercept we need to use the coordinates given. \((2,7)\rightarrow(x,y)\) We would just plug in those values for x and y in the equation above. \(\displaystyle (7)=\frac{2}{3}(2)+b\) Solve for b, \(\displaystyle b=\frac{17}{3}\) Bam, there we are! Make sense?

OpenStudy (anonymous):

OMG you saved me thank you!!!!

OpenStudy (austinl):

No problem!